# StreamDiT: Real-Time Streaming Text-to-Video Generation

### v3

@@ Abstract
denotes equal contributionsThis work was done while Akio was an intern at Meta from May to November 2024.
Recently, great progress has been achieved in text-to-video (T2V) generation by scaling transformer-based diffusion models to billions of parameters, which can generate high-quality videos. However, existing models typically produce only short clips offline, restricting their use cases in interactive and real-time applications. This paper addresses these challenges by proposing StreamDiT, a streaming video generation model. StreamDiT training is based on flow matching by adding a moving buffer. We design mixed training with different partitioning schemes of buffered frames to boost both content consistency and visual quality. StreamDiT modeling is based on adaLN DiT with varying time embedding and window attention. To practice the proposed method, we train a StreamDiT model with 4B parameters. In addition, we propose a multistep distillation method tailored for StreamDiT. Sampling distillation is performed in each segment of a chosen partitioning scheme. After distillation, the total number of function evaluations (NFEs) is reduced to the number of chunks in a buffer. Finally, our distilled model reaches real-time performance at 16 FPS on one GPU, which can generate video streams at 512p resolution. We evaluate our method through both quantitative metrics and human evaluation. Our model enables real-time applications, e.g. streaming generation, interactive generation, and video-to-video. We provide video results and more examples in our project website: https://cumulo-autumn.github.io/StreamDiT/.
@@ 1 Introduction
Video generation is a fundamental and trending problem in computer vision, with rapid progress in recent years. Diffusion models have been adopted for text-to-video (T2V) generation, and demonstrated capabilities of generating high-quality videos, with a potential as world simulators. The success was built upon the scalability of transformer-based architectures and huge compute resources to train billions of parameters. Despite achievements, many existing models are designed to generate short video clips, due to the high cost of scaling. As a result, generating long videos with low latency remains an extreme challenge [li2024surveylongvideogeneration], which is demanded in real-time and interactive applications.
Modern T2V models [moviegen, hunyuan, step-t2v] are based on Diffusion Transformers (DiTs) [Peebles2022DiT]. They use auto-encoders to compress a video clip to a 3D latent. The auto-encoders are based on CNN layers with uniform compression over spatial and temporal domains. Therefore, the latent frames are separable. The models predict all frames together by gradually removing noise from a latent. Increasing video length is expensive due to the quadratic complexity of transformers. To make diffusion models capable of generating long videos, some approaches [xie2024progressiveautoregressivevideodiffusion, yin2025slowbidirectionalfastautoregressive] attempted to combine autoregression and diffusion, by either adding progressive noise to latent frames or causal masking in the final model. In an autoregressive manner, a latent frame is predicted from previous ones. We argue that recent high-quality video generation models benefit from full attention of all tokens in a sequence, which ensures high consistency in video content. Therefore, a method that preserves token communication is desired for enabling extendable video generation.
@@
An inspiration comes from training-free methods for streaming generation, e.g. StreamDiffusion [kodaira2023streamdiffusionpipelinelevelsolutionrealtime] and FIFO-Diffusion [Kim2024FIFO]. StreamDiffusion utilizes pre-trained image diffusion models to edit an input video stream via batch denoising. It suffers from a lack of frame-to-frame consistency, leading to visual artifacts when generating continuous video streams. FIFO-Diffusion enqueues a sequence of frames with different noise levels. It pops up a clear frame after one denoising step and enqueues a noise frame in the queue. Although they do not require training, the quality is limited. Another missing effort is sampling distillation, which is crucial for real-time applications. Since the queue is updated at every denoising step, popular distillation methods e.g. guided distillation [guideddistillation] and consistency distillation [lcm] cannot be applied directly.
@@
To address the challenges, we propose a complete solution for streaming video generation, named StreamDiT, which includes training and modeling. For training, we modify the flow matching [lipman2023flowmatching] by introducing a moving buffer of frames. Within a buffer, we adopt full attention so that our method can be seamlessly applied to modern T2V models. At inference, the buffer is moving along the frame dimension, generating clean frames sequentially. Similar to previous work [xie2024progressiveautoregressivevideodiffusion, Kim2024FIFO], we allow buffered frames with different noise levels. We design a unified partitioning of latent frames with different chunk sizes and micro steps. We reveal that uniform noise used in original flow matching [lipman2023flowmatching] and diagonal noise from previous work [Kim2024FIFO, xie2024progressiveautoregressivevideodiffusion] are special cases of our partitioning. With mixed training of different schemes, we can enhance video consistency and avoid overfitting on a specific scheme.
@@
For modeling, we design our StreamDiT model architecture to fit its training process while being efficient for real-time applications. We modify the adaLN DiT [Peebles2022DiT] by adding varying time embedding and replacing full attention with window attention [swintransformer]. We first train our StreamDiT as a basic T2V model, and then adapt it to streaming video generation. To achieve real-time applications, we also design a multistep distillation tailored for StreamDiT with a chosen partitioning scheme. With that, we distill our model to 8 sampling steps without classifier-free guidance (CFG). Finally, our model can generate video frames at 16 FPS on one H100 GPU, achieving real-time performance.  Fig. 1 shows video results of our real-time streaming generation.
@@
  • We propose StreamDiT training for streaming video generation, which is based on flow matching by introducing a moving buffer. We design generalized partitioning of buffered frames that covers uniform noise, diagonal noise, and others in between. We also point out a requirement for generation models such that time embedding needs to be separable in the frame dimension.
@@
  • We build a real-time solution using multistep distillation tailored for StreamDiT. Choosing a partitioning scheme that balances inference speed and quality, we distill micro steps to one for each segment. With all the efforts combined, our distilled model achieves real-time performance with 16 FPS on one GPU. In addition, we show some results to highlight potential real-time applications.
@@ 2.1 Text-to-Video Generation
Arising with the development of text-to-image generation [dai2023emu], video generation bursts by introducing temporal modules. CogVideo [hong2022cogvideo] adds temporal attention modules into a pre-trained autoregressive text-to-image model CogView2 [ding2022cogview2]. Video Diffusion Models (VDM) [ho2022vdm] utilize a space-time factorized U-Net with joint image and video data training. Imagen Video [ho2022imagenvideo] improves VDM with cascaded diffusion models for high-resolution generation. Lumiere [bartal2024lumiere] introduces a Space-Time U-Net architecture that generates the entire temporal duration of the video at once. Noticing the high computation of generation in pixel domain, researchers explored video generation in latent domain, e.g[zhou2023magicvideo, he2023latentvideodiffusionmodels, blattmann2023alignlatents, guo2023animatediff, chen2024videocrafter2]. Yet, these methods are based on UNet structures with temporal layers added to T2I models. As revealed by the work of DiT [Peebles2022DiT], transformers show better scalability than UNets in T2I generation. VDT [Lu2024VDT] proposes to use separated spatial and temporal attentions in a transformer block. GenTron [chen2023gentron] introduces a family of transformer-based diffusion models for image and video generation. Snap Video [menapace2024snapvideo] proposes a transformer architecture with joint spatiotemporal blocks, which is trained faster than U-Nets. ViD-GPT [gao2024vidgpt] uses causal attention in temporal domain and frames as prompts, following the design of GPT in LLMs. Due to the limitations of compute, these models are trained at small or medium scales, up to a few of billion parameters, generating videos of a few seconds. In the latest work of MovieGen [moviegen], the DiT model has been scaled to 30B parameters and can generate videos up to 16 seconds, with high motion and aesthetic qualities.
@@
Recently, the open-source community is also growing fast. OpenSora [opensora] initiated releases of small-sized T2V models with about 1B parameters. They use factorized transformers for spatial and temporal domains. Hunyuan [hunyuan] video model has 13B parameters. It uses MMDiT [esser2024sd3] with concatenated text and video tokens in self-attention. For auto-encoder, it adopts causal 3D convolution [yu2024languagemodelbeatsdiffusion] in its VAE. Step-Video-T2V [step-t2v] with 30B parameters is by far the largest open-source model for video generation, which reduces the gap of model size between open-source and closed-source with similar quality.
2.2 Long Video Generation
@@
The literature of long video generation can be categorized as training and training-free methods. Training new models or new modules injected to existing models requires large compute resources. NUWA-XL [yin2023nuwaxl] proposes a diffusion over diffusion architecture to generate long videos in a coarse-to-fine manner. StreamingT2V [henschel2024streamingt2v] proposes an autoregressive approach for long video generation, using a short-term and long-term memory block of previous chunks. In [xie2024progressiveautoregressivevideodiffusion], a similar method was proposed by assigning latent frames with progressively increasing noise levels. VideoTetris [tian2024videotetris] trains a ControlNet branch of condition frames for long video generation. A recent work by Loong [wang2024loong] proposes an autoregressive LLM-based approach for video generation. The model is trained on 10-second videos and can be extended to one-minute long. Yet, the current quality of autoregressive models is not comparable with diffusion models.
@@
For training-free methods, Gen-L-Video [wang2023genlv] discovers that the denoising path of a long video can be approximated by joint denoising of overlapping short videos in the temporal domain. MimicMotion [zhang2024mimicmotion] extends MultiDiffusion [bar2023multidiffusion] to the temporal domain. It blends frames within overlap region progressively along the denoising process. Video-Infinity [tan2024videoinf] distributes a long-form video task to multiple GPUs, with dual-scope attention that modulates temporal self-attention to balance local and global contexts efficiently across the devices. FreeNoise [qiu2024freenoise] proposes a simple solution to generate consistent long videos by rescheduling a sequence of noises for long-range correlation and performing temporal attention over them by window-based function. Reuse-and-Diffuse [gu2023reusediffuse] proposes an iterative denoising for T2V, with staged guidance. It reuses intermediate denoising results from the previous clip to condition the current clip. FIFO-Diffusion [Kim2024FIFO] proposes an inference technique that iteratively performs diagonal denoising, which concurrently processes a series of consecutive frames with increasing noise levels in a queue.
@@ 2.3 Sampling Efficiency
Distillation of sampling steps has been widely adopted for accelerating diffusion models. Compared to advanced numerical solvers, sampling distillation demonstrates better quality with fewer number of sampling steps. Progressive distillation [salimans2021progressive] distills two denoising steps of a teacher model to one step of a student model. Guided distillation [guideddistillation] proposed to first distill two function evaluations of CFG to one, and then perform progressive distillation on sampling steps. Consistency distillation [song2023consistencymodels, lcm] maps any points on a diffusion trajectory to the same origin, and therefore reduces the number of steps needed for sampling. Multistep consistency models [heek2024multistepconsistencymodels] split the diffusion process to multiple segments, and consistency distillation can be applied at each segment. By allowing more budget on sampling steps, Multistep consistency distillation can achieve higher quality.
To further improve sampling efficiency, one-step approaches have been proposed, including UFOGen [ufogen] and DMD [dmd]. The idea is to directly match distributions without the Gaussian assumption. They require to hold multiple models (discriminator or score networks) during training, posing challenges for applying to large video models.
@@ 3.1 Buffered Flow Matching
Flow Matching: Our method is based on the Flow Matching (FM) [lipman2023flowmatching] framework for training. FM produces a sample from the target data distribution by progressively transforming a sample from an initial prior distribution, such as a Gaussian. During training, for a data sample denoted as $\mathbf{X}_{1}$, we sample a time step $t\in[0,1]$, and noise $\mathbf{X}_{0}\sim\mathcal{N}(0,1)$. These are then used to create a training sample $\mathbf{X}_{t}$. FM predicts the velocity $\mathbf{V}_{t}$ that moves the sample $\mathbf{X}_{t}$ in the direction of data sample $\mathbf{X}_{1}$.
A simple linear interpolation or the optimal transport (OT) path [lipman2023flowmatching] is used to construct $\mathbf{x}_{t}$, i.e.,
@@
Buffered Flow Matching: We consider streaming video generation as a sequence of (possibly latent) frames $f_{1},f_{2},\dots,f_{N}]$, and $N$ can be infinite. For a video diffusion model with frame buffer $B$, the clean data sample starting with frame $f_{i}$ is denoted as $\mathbf{X}_{1}^{i}=[f_{i},\dots,f_{i+B}]$. We allow different noise levels to the frames: $\tau=[\tau_{1},\dots,\tau_{B}]$ as a monotonically increasing sequence. Thus a training example can be constructed as
@@ 3.2 Partitioning Scheme
To facilitate flexible training and inference of StreamDiT, we design a unified partitioning of buffered frames. As illustrated in Fig. 2, the buffer is partitioned to $K$ reference frames and $N$ chunks. Each chunk has $c$ frames and $s$ micro denoising steps.
Clean Reference Frames: To enhance temporal consistency, we can optionally cache the last $K$ fully denoised frames at the beginning of the buffer. We refer to these clean frames as reference frames. They participate in the denoising step as input but are no longer denoised. The reference frames are updated in the same way as other frames when the buffer moves. Allowing optional reference frames matches the design of FIFO-Diffusion [Kim2024FIFO], which can be viewed as a special case. Since our method is trainable, we found that reference frames can be skipped for streaming video generation. Hence, we set $K=0$ in the rest of our work.
@@
Micro Step: As the total number of denoising steps is bounded by $N$, however, for better performance, additional denoising steps are usually necessary. A straightforward approach to address this limitation is to increase the size of the buffer, but this is limited by the maximum capacity of the model to denoise frames and would also increase latency for each frame in the buffer.
To overcome this, we introduce an additional dimension of the design called micro-denoising step, illustrated in Fig. 2. The core idea of micro step is to denoise stream chunks along the temporal axis while stagnating at a fixed spatial position for certain time. Let $s$ denote the denoising steps per micro step; then each stream chunk undergoes $s$ denoising steps before advancing to the next noise level and moving toward the output. This modification effectively extends the total number of denoising steps to $s\times N$ without increasing the buffer size.
@@
Mixed Training: As shown in Tab. 1, StreamDiT unifies original diffusion or FM with uniform noise and diagonal noise used by [Kim2024FIFO, xie2024progressiveautoregressivevideodiffusion]. The latter enables streaming generation but hurts consistency of generated content. To enhance consistency and avoid overfitting, we adopt mixed training with different schemes. This drives the model to learn generalized denoising with different noise levels, instead of memorizing fixed noise levels. It is worth noting that our mixed training covers diffusion and FM training. Therefore, our model can work as a standard T2V generation without streaming. This is also used for initializing our streaming generation.

### v4

@@ Abstract
denotes equal contributionsThis work was done while the first author was an intern at Meta.
Recently, great progress has been achieved in text-to-video (T2V) generation by scaling transformer-based diffusion models to billions of parameters. However, existing models typically produce only short clips offline, restricting their use cases in interactive and real-time applications. This paper addresses this challenge by proposing StreamDiT, an efficient trainable solution for streaming video generation. StreamDiT training modifies flow matching by adding a moving buffer with progressive denoising. We design mixed partitioning schemes of buffered frames to boost both content consistency and visual quality. StreamDiT modeling is based on adaLN DiT with varying time embedding and window attention. In addition, we propose a multistep distillation method tailored for StreamDiT. Sampling distillation is performed in each segment of a chosen partitioning scheme. After distillation, the total number of function evaluations (NFEs) is reduced to the number of chunks in a buffer. We apply StreamDiT to a 4B-parameter model, which reaches real-time performance for 512p resolution at 16 FPS on one GPU. We evaluate our method through both quantitative metrics and human evaluation. Our model enables real-time applications, e.g. streaming generation, interactive generation, and video-to-video editing. More video results are available at the project website: https://cumulo-autumn.github.io/StreamDiT/.
@@ 1 Introduction
Video generation is a trending problem in computer vision, with rapid progress in recent years. Diffusion models have been adopted for text-to-video (T2V) generation, and demonstrated capabilities of generating high-quality videos, with a potential as world simulators. The success was built upon the scalability of transformer-based architectures and huge compute resources to train billions of parameters. Despite achievements, many existing models are designed to generate short video clips, due to the increasing cost after scaling. As a result, generating long videos with low latency remains an extreme challenge [19], which is required in interactive and real-time applications.
Modern T2V models [30, 18, 25] are based on Diffusion Transformers (DiTs) [29] with bidirectional attention. Increasing video length is expensive due to the quadratic complexity of transformers with respect to sequence length. To make diffusion models capable of generating long videos, some approaches [39, 45, 14] attempted to combine autoregressive (AR) and diffusion models, by adding progressive noise or causal masking. In an AR setup, a generation frame can only see previous context frames in attention, posing a limitation on its quality. As shown in [20], replacing causal attention with bidirectional attention leads to a massive quality gain. Modern video generation models [48, 30, 18] are also based on bidirectional attention. On another axis, AR frame-by-frame prediction has low streaming latency. However, it still needs to wait for a full denoising process to obtain a clean output frame. Diffusion with bidirectional attention can have high throughput, but its latency is much lower than AR. These observations lead us to seek a solution with low latency, high throughput, and high quality for streaming video generation.
@@
We are inspired by non-trainable streaming generation e.g. StreamDiffusion [17] and FIFO-Diffusion [16]. StreamDiffusion utilizes pre-trained image diffusion models to edit an input video stream via batch denoising. It suffers from a lack of frame-to-frame consistency, leading to visual artifacts when generating continuous video streams. FIFO-Diffusion enqueues a sequence of frames with different noise levels. It pops up a clear frame after one denoising step and enqueues a noise frame in the queue. However, they do not come with any training solution, which limits their quality. Another missing effort is sampling distillation, which is crucial for real-time applications. Since the queue is updated at every denoising step, popular distillation methods e.g. step distillation [28] and consistency distillation [24] cannot be applied directly.
@@
To address the challenges, we propose a complete solution for streaming video generation including training, modeling, and distillation, named StreamDiT. For training, we modify the flow matching [21] by introducing a moving buffer of frames. Within a buffer, we adopt full attention on latent frames so that our method can be seamlessly applied to modern T2V models. At inference, the buffer is moving along the frame dimension, generating clean frames sequentially. Similar to previous work [39, 16], we allow buffered frames with different noise levels. We design a unified partitioning of latent frames with different chunk sizes and micro steps. The uniform noise [21] and the diagonal noise [16, 39] are special cases of our partitioning. With mixed training of different schemes, we can enhance video consistency and avoid overfitting on a specific scheme.
@@
For modeling, we design our StreamDiT model architecture to fit its training process while being efficient for real-time applications. We modify adaLN DiT [29] by adding varying time embedding and replacing full attention with window attention [22]. We first train our StreamDiT as a basic T2V model, and then adapt it to streaming video generation. To achieve real-time applications, we also design a multistep distillation tailored for StreamDiT with a chosen partitioning scheme. With that, we distill our model to 8 sampling steps without classifier-free guidance (CFG). Finally, our model can generate video frames at 16 FPS on one H100 GPU, achieving real-time performance.  Fig. 1 shows some video results of our real-time streaming generation.
@@
  • We propose StreamDiT training for streaming video generation, which is based on flow matching by introducing a moving buffer. We articulate a generalized partitioning of buffered frames that covers uniform noise, diagonal noise, and others in between. We also point out a requirement for generation models such that time embedding needs to be separable in the frame dimension.
@@
  • We build a real-time solution using multistep distillation tailored for StreamDiT. Choosing a partitioning scheme that balances inference speed and quality, we distill micro steps to one for each segment. With all the efforts combined, our distilled model achieves real-time performance with 16 FPS on one GPU. In addition, we show some results to highlight potential real-time applications using our model.
@@ 2.1 Text-to-Video Generation
Arising with the development of text-to-image generation (T2I) [5], video generation bursts by introducing temporal modules. Earlier work [49, 11, 2, 10, 3] was based on UNet structures with temporal layers added to T2I models. As a pivot, DiT [29] demonstrated that transformer architecture is more scalable than UNet architecture for diffusion models. VDT [23] proposes to use separated spatial and temporal attentions in a transformer block. GenTron [4] introduces a family of transformer-based diffusion models for image and video generation. Snap Video [26] proposes a transformer architecture with joint spatiotemporal blocks, which is trained faster than U-Nets. ViD-GPT [7] uses causal attention in temporal domain and frames as prompts, following the design of GPT in LLMs. Due to the limitations of compute, these models are trained at small or medium scales, up to a few of billion parameters, generating videos of a few seconds. In the latest work of MovieGen [30], the DiT model has been scaled to 30B parameters and can generate videos up to 16 seconds, with high motion and aesthetic qualities.
@@
In parallel, the open-source community is also growing fast. OpenSora [48] released a small T2V model with about 1B parameters, using factorized transformers for spatial and temporal domains. Hunyuan [18] video model reaches 13B parameters, which uses MMDiT [6] with concatenated text and video tokens in self-attention. For auto-encoder, it adopts causal 3D convolution [46] in its VAE. Step-Video-T2V [25] with 30B parameters is a large open-source model for video generation, which reduces the model size gap between open-source and closed-source.
2.2 Extensible Video Generation
@@
The literature of extensible long video generation can be roughly categorized as training and training-free methods. Training new models or new modules injected to existing models requires a large amount of compute resources. NUWA-XL [43] proposes a diffusion over diffusion architecture to generate long videos in a coarse-to-fine manner. StreamingT2V [13] proposes an autoregressive approach for long video generation, using a short-term memory block and long-term memory block of previous chunks. In [39], a similar method was proposed by assigning latent frames with progressively increasing noise levels. VideoTetris [36] trains a ControlNet branch of condition frames for long video generation. A recent work by Loong [38] proposes an autoregressive LLM-based approach for video generation. The model is trained on 10-second videos and can be extended to one-minute long. To reduce the training and inference gap of AR video generation, self-forcing [14] was proposed, which uses student predictions as previous frames in the context. To reduce the streaming latency, it generates one latent frame at a time. The current AR models are hard to compete with full-attention diffusion models on quality.
@@
For training-free methods, Gen-L-Video [37] discovers that the denoising path of a long video can be approximated by joint denoising of overlapping short videos in the temporal domain. MimicMotion [47] extends MultiDiffusion [1] to the temporal domain. It blends frames within overlap region progressively along the denoising process. Video-Infinity [34] distributes a long-form video task to multiple GPUs, with dual-scope attention that modulates temporal self-attention to balance local and global contexts efficiently across the devices. FreeNoise [31] proposes a simple solution to generate consistent long videos by rescheduling a sequence of noises for long-range correlation and performing temporal attention over them by window-based function. Reuse-and-Diffuse [9] proposes an iterative denoising for T2V, with staged guidance. It reuses intermediate denoising results from the previous clip to condition the current clip. FIFO-Diffusion [16] proposes an inference technique that iteratively performs diagonal denoising, which concurrently processes a series of consecutive frames with increasing noise levels in a queue.
@@ 2.3 Sampling Efficiency
Step distillation has been widely adopted for accelerating diffusion models. Compared to advanced numerical solvers, sampling distillation demonstrates better quality with fewer number of sampling steps. Progressive distillation [32] distills two denoising steps of a teacher model to one step of a student model. Guided distillation [28] proposed to first distill two function evaluations of CFG to one, and then perform progressive distillation on sampling steps. Consistency distillation [33, 24] maps any points on a diffusion trajectory to the same origin, and therefore reduces the number of steps needed for sampling. Multistep consistency models [12] splits the diffusion process to multiple segments, and consistency distillation can be applied at each segment. By allowing more budget on sampling steps, Multistep consistency distillation can achieve higher quality.
To further improve sampling efficiency, one-step approaches have been proposed, including UFOGen [41] and DMD [44]. The idea is to directly match distributions without the Gaussian assumption. They require to host auxiliary models (discriminator or fake score networks) during training, posing challenges for applying to large video models.
@@ 3.1 Buffered Flow Matching
Flow Matching: Our training is based on Flow Matching (FM) [21]. FM produces a sample from the target data distribution by progressively transforming a sample from an initial prior distribution, such as a Gaussian. For a data sample $\mathbf{X}_{1}$, FM samples a time step $t\in[0,1]$, and noise $\mathbf{X}_{0}\sim\mathcal{N}(0,1)$. These are used to create a training sample $\mathbf{X}_{t}$. FM predicts the velocity $\mathbf{V}_{t}$ that moves the sample $\mathbf{X}_{t}$ in the direction of data sample $\mathbf{X}_{1}$.
A simple linear interpolation or the optimal transport (OT) path [21] is used to construct $\mathbf{x}_{t}$, i.e.
@@
Buffered Flow Matching: We consider streaming video generation as a sequence of (possibly latent) frames $f_{1},f_{2},\dots,f_{N}]$, and $N$ can be infinite. For a video diffusion model with frame buffer $B$, the clean data sample starting with frame $f_{i}$ is denoted as $\mathbf{X}_{1}^{i}=[f_{i+1},\dots,f_{i+B}]$. We allow different noise levels to the frames: $\tau=[\tau_{1},\dots,\tau_{B}]$ as a monotonically increasing sequence. Thus a training example can be constructed as
@@ 3.2 Partitioning Scheme
To facilitate flexible training and inference of StreamDiT, we design a unified partitioning of buffered frames. As illustrated in Fig. 2, the buffer is partitioned to $K$ reference (context) frames and $N$ chunks. Each chunk has $c$ frames and $s$ micro denoising steps.
Context Reference Frames: To enhance temporal consistency, we can optionally cache the last $K$ fully denoised frames at the beginning of the buffer. We refer to these clean frames as reference frames. They participate in the denoising step as input but are no longer denoised. The reference frames are updated in the same way as other frames when the buffer moves. Allowing optional reference frames matches the design of FIFO-Diffusion [16], which can be viewed as a special case. Since our method is trainable, we found that reference frames can be skipped for streaming video generation. Hence, we set $K=0$ in the rest of our work.
@@
Micro Step: For better performance, additional denoising steps are usually helpful. A straightforward approach to address this limitation is to increase the size of the buffer, which will increase the serving cost, and therefore needs to be bounded by a computation budget.
To overcome this, we introduce an additional dimension of the design called micro-denoising step, illustrated in Fig. 2. The core idea of micro step is to denoise stream chunks along the temporal axis while stagnating at a fixed spatial position for a certain amount of time. Let $s$ denote the micro step; then each stream chunk undergoes $s$ denoising steps before advancing to the next noise level and moving toward the output. This modification effectively extends the total number of denoising steps to $s\times N$ without increasing the buffer size.
@@
Mixed Training: As shown in Tab. 1, StreamDiT unifies partion schemes from bidirectional attention with uniform noise to progressive diagonal noise [16, 39]. The latter enables streaming generation, but hurts consistency of generated content. To enhance consistency and avoid overfitting, we adopt mixed training with different schemes. This drives the model to learn generalized denoising with different noise levels, instead of memorizing fixed noise levels. It is worth noting that our mixed training covers diffusion and FM training. Therefore, our model can work as a standard T2V generation without streaming. This is also used to initialize our streaming generation.